OpenStudy (anonymous):

give solution 2sin(0/)+ sqrt3=0 interval 0,2pi

OpenStudy (freckles):

isolate the trig function just like you would isolate y in the equation 2y+sqrt(3)=0 (I'm going to use y instead so I can easily get you to use unit circle after)

OpenStudy (freckles):

sin(theta) is the y-coordinate on the unit circle

OpenStudy (anonymous):

so thats it @freckles

OpenStudy (freckles):

well did you find as the solutions?

OpenStudy (freckles):

but yes solved for y then find that y number on the unit circle

OpenStudy (anonymous):

is it y= sqrt 3/2 ? @freckles

OpenStudy (freckles):

y=-sqrt(3)/2 now where is y , -sqrt(3)/2 on the unit circle

OpenStudy (freckles):

hint: look at 3rd and fourth quadrants

OpenStudy (anonymous):

0,2 ?

OpenStudy (freckles):

are you looking at the unit circle you are looking to see where the y-coordinate is -sqrt(3)/2

OpenStudy (anonymous):

3/2

OpenStudy (freckles):

I will give you one location y coordinate is -sqrt(3)/2 when theta=4pi/3 you find the other location

OpenStudy (anonymous):

I dont really get it//

OpenStudy (freckles):

do you see the coordinates on the unit circle

OpenStudy (freckles):

the 2nd number is the y coordinate

OpenStudy (anonymous):

2y+sqrt(3)=0 @freckles